Solve the differential equation :
(x^2)y'' - y' - 4(x^3)y = 8(x^3)sin(x^2)
1
Expert's answer
2020-03-27T13:43:53-0400

We need to solve : "(x^2)y'' - y' - 4(x^3)y = 8(x^3)sin(x^2)"

Substituting "x^2=t \implies 2xdx=dt"

"\implies dy/dx=2xdy/dt" ---(1)

Differentiating this again, we get;

"d^2y\/dx^2=4xd^2y\/dt^2-(2\/x)(dy\/dt)" ---(2)

Using (1) and (2) in the given equation, we get

"4x^3(y''-y)=8x^3sint"

"\implies y''-y=2sint"

Solving this, we get;

"(D^2-1)y=2sint"

Complimentary function : "c_1e^t+c_2e^{-t}" (Since, roots of auxiliary equation are 1, -1)

Particular solution : "y=2sint\/(D^2-1)=2sint\/(-(1)^2-1)=-sint"

"=-sint"

Thus, final solution is "y=c_1e^t+c_2e^{-t}-sint"

Substituting "t=x^2" back, we get;

"y=c_1e^{x^2}+c_2e^{-x^2}-sin(x^2)"